OpenStudy (anonymous):

On a trip a motorist drove 10 miles at 20 miles per hour, 10 miles at 30 miles per hour, and 10 miles at 60 miles per hour. What fraction of her total driving time was spent driving 60 miles per hour?

OpenStudy (anonymous):

First I would divide into three sections of time. 1st: 10 miles at 20mph 2nd: 10 miles at 30mph 3rd: 10 miles at 60mph 1st: 10m / 20mph = 0.5h 2nd: 10m / 30mph = 0.33h 3rd: 10m / 60mph = 0.17h 0.5h + 0.33h + 0.17h = 1h Time going 60 / total time = 0.17/1 or 17/100 or 17% How's that?